# Dimensionality reduction techniques to support

### v1

@@ 4.1 Other case studies
Table 5: Anomaly detection on all assets, obtained by employing PCA. $N$ and $T$ are the numbers of investors and days in the data set. $K$ is the encoding dimension. $A$ is the set of investors identified as anomalous. $I$ is defined as $I=|A\cap A^{KM}|$ where $A_{KM}$ is set of investors identified as hard/soft by the method based on k-means. $I_{n}$ is defined as $I_{n}=|A_{n}\cap A^{KM}|$ where $A_{n}$ is the set of the first $n$ ranked anomalous investors.
Model $||X-\hat{X}||_{F}$ EVS $\bar{s^{*}}$ $\bar{s^{*}}_{anomalous}$ $\bar{s^{*}}_{normal}$ $M_{1}$ $M_{2}$
@@
Table 6: UBI: dimensionality reduction. EVS is the explained variance score and it is defined as $100\Big{(}1-\frac{Var(X-\hat{X})}{Var(X)}\Big{)}$; $\bar{s^{*}}$ is the mean anomaly score; $\bar{s^{*}}_{anomalous}$ is the mean anomaly score for investors who are identified as hard/soft by the method based on k-means; $\bar{s^{*}}_{normal}$ is the mean anomaly score for all investors except the ones identified as hard/soft by the method based on k-means; $M_{1}=\frac{\bar{s^{*}}_{anomalous}-\bar{s^{*}}_{normal}}{\bar{s^{*}}_{normal}}$; $M_{2}=\frac{\bar{s^{*}}_{anomalous}}{\bar{s^{*}}}$. The results related to the autoencoders are averaged over $10$ runs.
While we have extensively covered the case study related to the asset IMA, we now focus on the other PSEs shown in Table 1. Table 5 summarizes the main results obtained by using PCA. The overlapping with the results of the method based on k-means are analogous to what is obtained for IMA, except for MOLMED. For this asset, the small overlapping is due to the choice of the investigation period, thus highlighting the ability of our method based on a dimensionality reduction approach to be independent of the choice of the investigation period. Moreover, the value added by our new method to the insider trading detection task, is analogous to what is obtained for IMA.
Now, let us deepen the main results related to the asset UBI. We employ both PCA and autoencoders. In Table 6, a comparison between the reconstruction results obtained by employing different architectures is shown. A trend different from IMA can be observed. In this case, we need at least $3$ hidden layers to outperform PCA in the reconstruction of trading profiles and AE-3 shows the lowest MSE. Contrary to IMA, the architecture which leads to the greatest values of $M_{1}$ and $M_{2}$ is still AE-3. If we rely on this autoencoder and apply our anomaly detection method, the results we obtain are provided in Table 7, compared with the ones of PCA.
@@
Table 7: UBI: anomaly detection. The encoding dimension is $16$. $A$ is the set of investors identified as anomalous. $I$ is defined as $I=|A\cap A^{KM}|$ where $A_{KM}$ is set of investors identified as hard/soft by the method based on k-means. $I_{n}$ is defined as $I_{n}=|A_{n}\cap A^{KM}|$ where $A_{n}$ is the set of the first $n$ ranked anomalous investors.
If the first $500$ ranked by PCA are considered, our new method does not provide new information compared to the one based on k-means. This is in contrast with the autoencoder. The profiles detected by AE-3 and not by k-means are analogous to the one in the top left panel of Figure 4. On the other hand, the investors detected by AE-3 and not by PCA are hard according to the method based on k-means and analogous to the profile in Figure 17. The ability of autoencoders to capture this type of investors which are not detected by PCA, was already shown in the study related to IMA.
5 Conclusion
@@
List of abbreviations
  • AE: Autoencoder
  • CARRARO: Carraro S.p.A.
  • EVS: Explained Variance Score
  • IMA: Industria Macchine Automatiche S.p.A.
  • LAE: Linear Autoencoder
  • LSTM: Long Short-Term Memory
  • MOLMED: MolMed S.p.A.
  • MSE: Mean Squared Error
  • PANARIAGROUP: Panariagroup Industria Ceramiche S.p.A.
  • PCA: Principal Component Analysis
  • PSE: Price Sensitive Event
  • ReLU: Rectified Linear Unit
  • SVD: Singular Value Decomposition
  • UBI: UBI Banca - Unione di Banche Italiane S.p.A.
@@
Availability of data and materials
The data that support the findings of this study are from the Commissione Nazionale per le Società e la Borsa (CONSOB), that is the public authority responsible for regulating the Italian financial markets. Restrictions apply to the availability of these data, because of the severe privacy policy related to the data collected within the MiFIDII/MiFIR regime and are not publicly available.
Competing interests
The authors declare that they have no competing interests.
Funding
This paper is funded by the project “Percorso di formazione su data analysis, network science, artificial intelligence e big data per gli abusi di mercato”, a research collaboration between Scuola Normale Superiore and CONSOB.
Authors’ contributions
Adele Ravagnani and Fabrizio Lillo contributed to the conception and design of the work and the interpretation of the results. Adele Ravagnani developed the methodology, analyzed the data, and was the major contributor to writing the manuscript. All authors read and approved the final manuscript.
APPENDICES
@@ Appendix A The choice of $K$ in PCA
In Figure 7, the trading profile of an investor is compared with the reconstructed ones obtained by PCA for several values of $K$. It is evident that increasing the latent space dimension leads to an improvement in the reconstruction of the profile. Moreover, $K=16$ allows to obtain reconstruction errors which are comparable to the ones related to higher values of $K$. Indeed, the choice $K=16$ is such that $97\%$ of the data variance can be explained. However, as we illustrated in Subsection 2.2, giving the unsupervised nature of our problem and our complete lack of labels, a priori we do not know the best suited value of this parameter. This motivates us to perform an analysis to understand how different choices of $K$ could impact our results.
@@ Appendix B PCA on data in two different formats
Our starting data set is in the format $N\times T$ i.e. the trading days are the features which are subjected to compression. We could start with a data set $Y\in\mathbb{R}^{T,N}$ where the features are the investors. As we explained in Subsection 2.2, this would lead to a more time consuming and computationally expensive procedure since in our dataset $N\gg T$. However, it is legitimate to ask whether there is a difference in the results obtained with these two approaches.
@@
$x_{i}(t)\rightarrow\frac{x_{i}(t)-mean(x_{t})}{std(x_{t})}$ (7)
@@
$x_{i}(t)\rightarrow\frac{x_{i}(t)-mean(x_{i})}{std(x_{i})}.$
@@
$(XX^{T})(Xp)=\lambda(Xp)\iff Cov(Y)(Xp)=\lambda(Xp)$
where $Y=X^{T}$. Therefore, the eigenvalues of $Cov(X)$ and $Cov(Y)$ are the same while the eigenvectors are $p$ and $Xp$.
@@ Appendix C Relation between linear autoencoders and PCA
If we compare Equations 4 and 6, it is pretty evident they are analogous if the activation functions $g_{2}$ and $g_{1}$ are the identity functions and $W_{1}=W_{2}^{T}$. Indeed, as illustrated in [16], if the activation functions are linear, the autoencoder with $L_{2}$-regularization learn PCA’s principal directions.
@@ Appendix D Anomaly detection with autoencoders
In this section, figures concerning the results related to our method based on the employment of autoencoders and applied on the asset IMA, are provided. Explanations of these results are inserted in the main text.
@@ Appendix E Households versus firms
As we explained in Section 3, the data set we are provided with, contains information about each investor type. We consider two broad categories of investors: households, that include individual households and joint accounts of several households, and firms, that include investment firms and other legal entities. The goal of this section is to investigate whether it could be more advantageous to run our whole methodology separately for households and firms. This is motivated by the fact that, in principle, these two classes of investors have different behavior.
@@
Table 8: IMA and UBI. Composition of the two groups of investors in the latent space and in the anomaly score distribution.
E.1 PCA using all data: household-firm composition
@@
We consider the representation in the latent space obtained by PCA which uses all data. A clustering method (the k-means) for two groups is run and we check whether one of the group is mainly composed of households and the other one of firms. The groups’ composition is reported in Table 8. We perform a Fisher test with the null hypothesis that there is not association between groups in the latent space and investor types. The p-value turns out to be $4.3$e-$5$ so, we need to reject the null hypothesis: there is a relation between investor type and group. Analogously, we test whether a given investor type is over/under-expressed in a group, as in [24]. For this test, the null hypothesis is defined by assuming the random co-occurrence of a given investor type and her belonging to a given group. The hypergeometric distribution is used as a benchmark for randomness. It results that in group 1 (2), the investor type household (firm) is over-expressed and the investor type firm (household) is under-expressed.
However, given that our method relies on the computation of the reconstruction errors, we further investigate the household-firm composition in the anomaly score distribution obtained by running PCA using the whole dataset. We split investors in two categories: investors with anomaly score $s_{i}^{*}$ greater than or equal to the threshold $\epsilon_{\theta}$ (group A) and investors with anomaly score lower than the threshold (group B). The groups’ composition is shown in Table 8. Also in this case, the Fisher test points out that there is association between the investor type and the group in the anomaly score distribution. We test the over/under-expression of investor types in the two groups, as in [24]. It results that in group A (B), the investor type firm (household) is over-expressed and the investor type household (firm) is under-expressed. We can conclude that, basically, firms are associated with higher values of anomaly score.
@@
E.2 PCA using households’ and firms’ data separately
@@
E.3 Insider trading detection using households’ and firms’ data separately
Now, let us tackle our major goal of this section, that is investigating whether it could be more advantageous to run our whole methodology for insider trading detection, separately for households and firms. As we illustrated in the previous subsection, for IMA (UBI), firms (households) are associated with higher values of anomaly score and households (firms) with lower values of anomaly score. For IMA, this issue together with the small number of firms ($606+39$) imply that, if we perform PCA using only the data related to firms, the anomaly score distribution we obtain, does not show the bimodality we want to exploit in order to set the threshold $\epsilon_{\theta}$, which has a major role in the criterion of Equation 3. On the other hand, for UBI, the higher number of firms (in absolute value) allows to preserve the bimodality of the anomaly score distribution obtained by running PCA with only the data related to firms, as shown in Figure 20. Therefore, we focus on UBI for the subsequent analysis.
We apply our whole methodology to identify potential insiders, for households and firms separately: results are shown in Table 9. The method which uses only the data related to firms identifies $16$ more anomalous investors than the method which uses the whole dataset. Some of them, like the profile in Figure 21, could be interesting for our scope. However, if we focus on the first $500$ ranked potential insiders, there is no difference. On the other hand, if the methodology is run by using only the data related to households, a consistent number of potential insiders is not identified with respect to the method which uses all the data and, among the first $500$ ranked potential insiders, $59$ investors are not detected. These households are actually extremely suspicious since they are all just active in the investigation period with a net buying position, similarly to the profile in Figure 21.
The difference between the results obtained by using all data and the data only related to households, could be surprising: in subsection E.2, we observed that the difference between the principal components and the eigenvalues obtained in the two cases is negligible. However, it is important to remember that in the criterion of Equation 3, also the times corresponding to the largest reconstruction errors $t_{i}^{*}$ have a role and in fact, using the data only related to households, causes a change in the $t_{i}^{*}$ histogram.
@@ E.3 Insider trading detection using households’ and firms’ data separately
Table 9: UBI. $A_{c}$ with $c=\{\text{households},\ \text{firms}\}$ is the set of potential insiders obtained by using only the data related to investors of type $c$. $A^{all}$ is the set of potential insiders obtained by using all data. $C$ is the set of households/firms in the data set. $A^{all}_{500}$ and $A_{500}^{c}$ are the set of the first $500$ ranked potential insiders obtained by using all data or only the data of investors of type $c$ respectively.

### v2

@@ IV-D Other case studies
TABLE VI: UBI: dimensionality reduction. The results related to the autoencoders are averaged over $10$ runs.
Model $||X-\hat{X}||_{F}$ EVS $\bar{s^{*}}$ $\bar{s^{*}}_{anomalous}$ $\bar{s^{*}}_{normal}$ $\text{M}_{1}$ $\text{M}_{2}$
@@
While we have extensively covered the case study related to the asset IMA, we now focus on the other PSEs shown in Table I. Table V summarizes the main results obtained by using PCA. The overlapping with the results of the method based on k-means are analogous to what is obtained for IMA, except for MOLMED. For this asset, the small overlapping is due to the choice of the investigation period, thus highlighting the ability of our method based on a dimensionality reduction approach to be independent of the choice of the investigation period. Moreover, the value added by our new method to the insider trading detection task, is analogous to what is obtained for IMA.
Now, let us deepen the main results related to the asset UBI. We employ both PCA and autoencoders. In Table VI, a comparison between the reconstruction results obtained by employing different architectures is shown. A trend different from IMA can be observed. In this case, we need at least $3$ hidden layers to outperform PCA in the reconstruction of trading profiles and AE-3 shows the lowest MSE. Contrary to IMA, the architecture which leads to the greatest values of $\text{M}_{1}$ and $\text{M}_{2}$ is still AE-3. If we rely on this autoencoder and apply our anomaly detection method, the results we obtain are provided in Table VII, compared with the ones of PCA.
TABLE VII: UBI: anomaly detection. The encoding dimension is $16$. $A$ is the set of investors identified as anomalous. $I$ is defined as $I=|A\cap A^{KM}|$ where $A_{KM}$ is set of investors identified as hard/soft by the method based on k-means. $I_{n}$ is defined as $I_{n}=|A_{n}\cap A^{KM}|$ where $A_{n}$ is the set of the first $n$ ranked anomalous investors.
@@
If the first $500$ ranked by PCA are considered, our new method does not provide new information compared to the one based on k-means. This is in contrast with the autoencoder. The profiles detected by AE-3 and not by k-means are analogous to the one in the top left panel of Figure 4. On the other hand, the investors detected by AE-3 and not by PCA are hard according to the method based on k-means and analogous to the profile in Figure 17 (Appendix D). The ability of autoencoders to capture this type of investors which are not detected by PCA, was already shown in the study related to IMA.
V Conclusion
@@
This paper represents the personal opinions of the authors and does not bind the membership organization in any way.
Restrictions apply to the availability of the data, because of the severe privacy policy related to the data collected within the MiFIDII/MiFIR regime and are not publicly available.
@@ Appendix A The choice of $K$ in PCA
In Figure 7, the trading profile of an investor is compared with the reconstructed ones obtained by PCA for several values of $K$. It is evident that increasing the latent space dimension leads to an improvement in the reconstruction of the profile. Moreover, $K=16$ allows to obtain reconstruction errors which are comparable to the ones related to higher values of $K$. Indeed, the choice $K=16$ is such that $97\%$ of the data variance can be explained. However, as we illustrated in Subsection II-B, giving the unsupervised nature of our problem and our complete lack of labels, a priori we do not know the best suited value of this parameter. This motivates us to perform an analysis to understand how different choices of $K$ could impact our results.
@@ Appendix B PCA on data in two different formats
Our starting data set is in the format $N\times T$ i.e. the trading days are the features which are subjected to compression. We could start with a data set $Y\in\mathbb{R}^{T,N}$ where the features are the investors. As we explained in Subsection II-B, this would lead to a more time consuming and computationally expensive procedure since in our dataset $N\gg T$. However, it is legitimate to ask whether there is a difference in the results obtained with these two approaches.
@@
$x_{i}(t)\rightarrow\frac{x_{i}(t)-\text{mean}(x_{t})}{\text{std}(x_{t})}$ (7)
@@
$x_{i}(t)\rightarrow\frac{x_{i}(t)-\text{mean}(x_{i})}{\text{std}(x_{i})}.$
@@
$(XX^{T})(Xp)=\lambda(Xp)\iff\text{Cov}(Y)(Xp)=\lambda(Xp)$
where $Y=X^{T}$. Therefore, the eigenvalues of $\text{Cov}(X)$ and $\text{Cov}(Y)$ are the same while the eigenvectors are $p$ and $Xp$.
@@ Appendix C Relation between linear autoencoders and PCA
If we compare Equations 4 and 6, it is pretty evident they are analogous if the activation functions $g_{2}$ and $g_{1}$ are the identity functions and $W_{1}=W_{2}^{T}$. Indeed, as illustrated in [32], if the activation functions are linear, the autoencoder with $L_{2}$-regularization learns PCA’s principal directions.
@@ Appendix D Anomaly detection with autoencoders
In this appendix, figures concerning the results related to our method based on the employment of autoencoders and applied on the asset IMA, are provided. Explanations of these results are inserted in the main text.
@@ Appendix E Households versus firms
As we explained in Section III, the data set we are provided with, contains information about each investor type. We consider two broad categories of investors: households, that include individual households and joint accounts of several households, and firms, that include investment firms and other legal entities. The goal of this appendix is to investigate whether it could be more advantageous to run our whole methodology separately for households and firms. This is motivated by the fact that, in principle, these two classes of investors have different behavior.
TABLE VIII: IMA and UBI. Composition of the two groups of investors in the latent space and in the anomaly score distribution.
@@
E-A PCA using all data: household-firm composition
@@
We consider the representation in the latent space obtained by PCA which uses all data. A clustering method (the k-means) for two groups is run and we check whether one of the group is mainly composed of households and the other one of firms. The groups’ composition is reported in Table VIII. We perform a Fisher test with the null hypothesis that there is not association between groups in the latent space and investor types. The p-value turns out to be $4.3$e-$5$ so, we need to reject the null hypothesis: there is a relation between investor type and group. Analogously, we test whether a given investor type is over/under-expressed in a group, as in [39]. For this test, the null hypothesis is defined by assuming the random co-occurrence of a given investor type and her belonging to a given group. The hypergeometric distribution is used as a benchmark for randomness. It results that in group 1 (2), the investor type household (firm) is over-expressed and the investor type firm (household) is under-expressed.
However, given that our method relies on the computation of the reconstruction errors, we further investigate the household-firm composition in the anomaly score distribution obtained by running PCA using the whole dataset. We split investors in two categories: investors with anomaly score $s_{i}^{*}$ greater than or equal to the threshold $\epsilon_{\theta}$ (group A) and investors with anomaly score lower than the threshold (group B). The groups’ composition is shown in Table VIII. Also in this case, the Fisher test points out that there is association between the investor type and the group in the anomaly score distribution. We test the over/under-expression of investor types in the two groups, as in [39]. It results that in group A (B), the investor type firm (household) is over-expressed and the investor type household (firm) is under-expressed. We can conclude that, basically, firms are associated with higher values of anomaly score.
@@
E-B PCA using households’ and firms’ data separately
@@
E-C Insider trading detection using households’ and firms’ data separately
Now, let us tackle our major goal of this appendix, that is investigating whether it could be more advantageous to run our whole methodology for insider trading detection, separately for households and firms. As we illustrated in the previous paragraph, for IMA (UBI), firms (households) are associated with higher values of anomaly score and households (firms) with lower values of anomaly score. For IMA, this issue together with the small number of firms ($606+39$) imply that, if we perform PCA using only the data related to firms, the anomaly score distribution we obtain, does not show the bimodality we want to exploit in order to set the threshold $\epsilon_{\theta}$, which has a major role in the criterion of Equation 3. On the other hand, for UBI, the higher number of firms (in absolute value) allows to preserve the bimodality of the anomaly score distribution obtained by running PCA with only the data related to firms, as shown in Figure 20. Therefore, we focus on UBI for the subsequent analysis.
We apply our whole methodology to identify potential insiders, for households and firms separately: results are shown in Table IX. The method which uses only the data related to firms identifies $16$ more anomalous investors than the method which uses the whole dataset. Some of them, like the profile in Figure 21, could be interesting for our scope. However, if we focus on the first $500$ ranked potential insiders, there is no difference. On the other hand, if the methodology is run by using only the data related to households, a consistent number of potential insiders is not identified with respect to the method which uses all the data and, among the first $500$ ranked potential insiders, $59$ investors are not detected. These households are actually extremely suspicious since they are all just active in the investigation period with a net buying position, similarly to the profile in Figure 21.
The difference between the results obtained by using all data and the data only related to households, could be surprising: in Subsection E-B, we observed that the difference between the principal components and the eigenvalues obtained in the two cases is negligible. However, it is important to remember that in the criterion of Equation 3, also the times corresponding to the largest reconstruction errors $t_{i}^{*}$ have a role and in fact, using the data only related to households, causes a change in the $t_{i}^{*}$ histogram.
TABLE IX: UBI. $A_{c}$ with $c=\{\text{households},\ \text{firms}\}$ is the set of potential insiders obtained by using only the data related to investors of type $c$. $A^{all}$ is the set of potential insiders obtained by using all data. $C$ is the set of households/firms in the data set. $A^{all}_{500}$ and $A_{500}^{c}$ are the set of the first $500$ ranked potential insiders obtained by using all data or only the data of investors of type $c$ respectively.